Hi there I need help figuring out what's to do for the following question.

The path of a ball is modeled by the quadratic function h(t)=-5(t-2)^2+23, where height, h(t), is in meters and time, t, is in seconds.
1) When will the ball reach a height of 18 m?

Answer by Alan3354(69443)

h(t)=-5(t-2)^2+23 = 18
-5(t-2)^2 + 23 = 18
Solve for t
There are 2 times, going up, and falling.